# BABILong: Testing the Limits of LLMs with Long Context Reasoning-in-a-Haystack

### v1

@@ Abstract
In recent years, the input context sizes of large language models (LLMs) have increased dramatically. However, existing evaluation methods have not kept pace, failing to comprehensively assess the efficiency of models in handling long contexts. To bridge this gap, we introduce the BABILong benchmark, designed to test language models’ ability to reason across facts distributed in extremely long documents. BABILong includes a diverse set of 20 reasoning tasks, including fact chaining, simple induction, deduction, counting, and handling lists/sets. These tasks are challenging on their own, and even more demanding when the required facts are scattered across long natural text. Our evaluations show that popular LLMs effectively utilize only 10-20% of the context and their performance declines sharply with increased reasoning complexity. Among alternatives to in-context reasoning, Retrieval-Augmented Generation methods achieve a modest 60% accuracy on single-fact question answering, independent of context length. Among context extension methods, the highest performance is demonstrated by recurrent memory transformers, enabling the processing of lengths up to 11 million tokens. The BABILong benchmark is extendable to any length to support the evaluation of new upcoming models with increased capabilities, and we provide splits up to 1 million token lengths.
@@ 1 Introduction
Today, large language models (LLMs) and neural architectures are continually evolving and achieving remarkable improvements, particularly in their ability to handle longer contexts (OpenAI, 2023; Reid et al., 2024; Anthropic, 2024). The ability of these models to process and generate text based on rich contextual information is crucial for several reasons. For example, longer contexts provide more information for the model to condition its outputs, leading to more accurate, contextually relevant, and up-to-date responses. Furthermore, long-context capabilities can enhance in-context learning by providing more in-context examples, instructions to follow, or example trajectories in context of reinforcement learning (Chevalier et al., 2023; Agarwal et al., 2024; Lee et al., 2024).
@@
Despite these advances in models capabilities, the benchmarks used to evaluate them have not kept pace. For example, current benchmarks such as Longbench (Bai et al., 2023) and L-Eval An et al. (2023) scale only up to 40,000 tokens, while models are capable of hundreds of thousands and millions of tokens (OpenAI, 2023; Bulatov et al., 2024; Gu & Dao, 2023; Anthropic, 2024; Reid et al., 2024; Liu et al., 2024a).
@@
Figure 1: a) Generation of BABILong dataset. Facts relevant for the question are hidden inside a larger background texts from PG19. b) Recurrent transformers answer questions about facts from very long texts when retrieval augmented generation fails. Common RAG method fails to answer questions because order of facts matters. GPT-4 LLM effectively uses only about 10% of the full 128K window. Small LMs, RMT with GPT-2 (137M) and Mamba (130M) fine-tuned for the task are able to solve it, with recurrent memory transformer scoring well up to record 11 111 111 tokens.
@@
Creating natural and comprehensive long-context benchmarks that are human labeled is very challenging. As a consequence, synthetic benchmarks focusing on variations of "needle-in-a-haystack" tasks have become increasingly common (Zhang et al., 2024b; Liu et al., 2024a; Song et al., 2024b; Hsieh et al., 2024). One widely used needle-in-a-haystack task involves finding specific "needles with magic numbers" in a haystack of Paul Graham’s essays [^1] https://github.com/gkamradt/LLMTest_NeedleInAHaystack. However, the widespread use of this approach has highlighted its limitations - it is overly simplistic, and novel long context models often achieve perfect performance, as usually demonstrated by fully green heatmaps (Reid et al., 2024; Cohere, 2024; Liu et al., 2024a; Wang et al., 2024c). This shows that while it serves well as a basic verification tool, it is not a rigorous benchmark that can effectively challenge and differentiate advanced long-context models.
@@
To bridge this gap, we introduce the BABILong benchmark, designed to test language models’ ability to reason across facts distributed in extremely long documents. BABILong includes a diverse set of 20 reasoning tasks, including fact chaining, simple induction, deduction, counting, and handling lists/sets, that were designed as prerequisites for any system that aims to be capable of conversing with a human (Weston et al., 2016). As a source of long natural documents we use books from PG19 corpora (Rae et al., 2020). In this way, BABILong allows the construction of tasks of almost arbitrary length, in order to adapt them to the evaluation of new, more powerful models in an extensible and controllable way. We provide sets of predefined lengths with splits up to 1 million tokens, and we evaluate models on samples with up to 11 million tokens.
@@
We find that popular LLMs effectively use only 10-20% of the context, with performance declining sharply as length and task complexity increase. Retrieval-Augmented Generation methods achieve a modest 60% accuracy in answering single-fact questions, regardless of context length. Among other methods, Mamba and Recurrent Memory Transformers (RMT) show the highest performance, with RMT capable of processing lengths up to 11 million tokens.
@@
2. We evaluate over 20 recent long-input language models with various sizes, architectures, and context extension methods on BABILong.
@@
4. We demonstrate successful in domain single fact question answering with the recurrent memory transformer on input texts up to 11 million tokens, setting a new record for the sequence size processed by a single model, extending the known capabilities of neural networks.
The BABILong benchmark data and code for evaluation are available on GitHub  [^2] https://github.com/booydar/babilong.
@@ 2 The BABILong Benchmark for Long Context Processing
The fundamental concept behind the Benchmark for Artificial Intelligence for Long-context evaluation is to extend the length of existing tasks to evaluate the ability of generative models to efficiently handle long contexts. Solving tasks with a long context size requires the model to distinguish important information from large amounts of irrelevant details. To simulate this behavior we "hide" the sentences of the original task between the sentences of irrelevant text that is drawn from another closely related distribution (see Figure 1a). Examples are constructed by gradually adding new sentences from the background dataset in their natural order until the augmented sample reaches the desired length. This way, we are not bound by the length of the original task itself, making it possible to assess even the longest available models with context sizes up to millions of tokens. For background text we use books from the PG19 dataset  (Rae et al., 2020) due to the substantial book lengths and naturally occurring long contexts. The model is required first to distinguish the sentences related to the original task, then memorize and subsequently utilize them to generate the correct solution.
@@ 3 Benchmarking Results
Table 2: BABILong is a challenging benchmark for current long-context models. Even models that claim to support 128K tokens experience degradation beyond 10% of their input capacity. RAG methods do not help, while fine-tuning of small scale models (RMT 137M and Mamba 130M) shows that the tasks are solvable. Values represent average accuracy over QA1-QA5 tasks from BABILong.
To maximize value for the research community, we have included models with the highest number of monthly downloads from the Hugging Face platform in our evaluation such as LLama-3 (AI@Meta, 2024); 32k-64k – Mistral (Jiang et al., 2023), Mixtral (Jiang et al., 2024); 128k – ChatGLM3 (Du et al., 2022), Phi-3 (Abdin et al., 2024), Command-R (Cohere, 2024); 200k – Yi (Young et al., 2024); including long-context fine-tuning: LongChat (Li et al., 2023a), LLama-2-7b-32k  [^3] https://huggingface.co/togethercomputer/Llama-2-7B-32K-Instruct, LongAlpaca (Chen et al., 2023); long-context adaptation methods: Yarnv2 Mistral (Peng et al., 2023b), Mistral and LLama-2 with Activation Beacons (Zhang et al., 2024a). As a reference, we included GPT-4 (gpt-4-0125-preview), currently the most powerful model available. Retrieval-augmented generation was also tested, as it represents a common solution for long document QA. As alternatives to traditional architectures, we considered the Mamba (Gu & Dao, 2023), Jamba (Lieber et al., 2024) and Recurrent Memory Transformer (RMT) (Bulatov et al., 2022). Summary of evaluation results is presented in the Table 2. The table reports average accuracy of models on the first 5 tasks (QA1-QA5) of BABILong for different context sizes. For evaluation details for each task see Table 3 and Appendix C.
@@ 3.1 Evaluation of Effective Context Size
Our benchmarking results show that current LLMs do not efficiently use their full context (Fig.2). Only 15 out of 24 tested LLMs were able to correctly answer 85% or more of the questions for any of QA1-QA3 tasks in a baseline setting without any background distractor text. Even for the the simplest task involving a single supporting fact (QA1), the best models are able to use only up to 4K tokens efficiently, except for GPT-4, which performs well till 16K. The range of full context utilization varies from 5% to 25%. When two supporting facts are required for an answer, only GPT-4 can solve the task without background text. When facts are embedded within texts, all tested LLMs fall below 85% performance(Fig.2QA2). The task with three supporting facts proves extremely challenging to current LLMs, with the best scores falling below 70% (Fig.2QA3).
Going deeper in performance of specific models presented in the Table 2 we found the following. Yarn fails to extend to longer contexts despite showing stable results in long-context language modeling (Peng et al., 2023b). LongChat, LongAlpaca, and both LLama2-7B-32K and LLama2-7B-32K-instruct models, even when fine-tuned on 32K lengths, failed to perform well on 32K tokens. Activation Beacon performed better than Yarn context extension method for Mistral 7B, but still achieved low results (< 40%) on 32K contexts. In contrast, Mistral and Mixtral, trained on lengths up to 32K, performed well on these lengths. Yi-9B-200k, trained on sequences up to 200K, shows less than 30% on 64K tokens and more. Yi-34B-200k shows very promising and stable results on lengths up to 64K, but unfortunately we were not able to run it on 128K tokens. Phi-3-Mini drops significantly from 64K to 128K, reaching less than 10%, while Phi-3-Medium maintains 30% at 128K. Jamba-v1 and Phi-3-Mini show close results, but Jamba-v1 does not have drop at 128K and shows 34% on this length. Command-R and Phi-3-Medium are the most robust to longer contexts, but start to lose performance more sharply at 128K. Phi-3-Medium and Command-R show results very close to GPT-4 at 32K+ contexts.
@@ 3.3 Fine-Tuning Models on BABILong
We performed fine-tuning experiments with GPT-3.5-Turbo, Mistral-7B-Instruct-v0.2, RMT with GPT-2 (137M) backbone, and Mamba (130M) models. Fine-tuning results are in Figure 3b and Appendix I Figure 8.
RMT with a GPT-2 (Radford et al., 2019) backbone model is trained on each task individually with a segment size of 512 and memory size of 16. Train and evaluation splits of each task contain 10000 and 1000 samples, respectively, with a number of facts in each sample ranging from 2 to 320 depending on the task. A curriculum strategy is employed, initiating training on short sequences that fit in a single segment and then gradually introducing longer sequences once the training converges. During each curriculum stage $n$, sequence lengths are randomly sampled from $1$ to $n$ segments. We provide details of training and evaluation procedures in Appendix E.
RMT model trained on 32 segments totalling in 16K tokens demonstrates strong performance on this length. Notably, RMT outperforms GPT-4 significantly, underscoring the efficiency of memory mechanism in processing long context. Even more importantly, the power of recurrent models extends to sequences longer than the training size. RMT shows consistent performance on longer sequences, up to 128k tokens, with only a marginal quality degradation. Surprisingly, with context sizes scaling to 1 million, 10 million tokens, and even 11.1 million tokens, which is over 600 times of the training length, the recurrent model persistently outperform the larger counterparts utilizing RAG.
Finetuned recurrent models, Mamba and RMT perform equally well on QA1, however due to the technical limitations of the Mamba implementation, the inference beyond 128k was extremely slow, which makes it nearly impossible to process longer sequences. RMT greatly outperforms retrieval-augmented models and is able to process sequences up to 10M tokens much faster than Mamba. However, Mamba has an edge in complex tasks such as remembering a large number of facts in QA3.
We evaluated GPT-3.5 and Mistral-7B models fine-tuned with 1000 samples from QA1 for 3 epochs. The evaluation results are shown in Appendix I Figure 8. Fine-tuning dramatically improves performance for longer contexts making scores uniform across all input sizes. Still, these results are behind of fine-tuned Mamba and RMT.

### v2

@@ Abstract
In recent years, the input context sizes of large language models (LLMs) have increased dramatically. However, existing evaluation methods have not kept pace, failing to comprehensively assess the efficiency of models in handling long contexts. To bridge this gap, we introduce the BABILong benchmark, designed to test language models’ ability to reason across facts distributed in extremely long documents. BABILong includes a diverse set of 20 reasoning tasks, including fact chaining, simple induction, deduction, counting, and handling lists/sets. These tasks are challenging on their own, and even more demanding when the required facts are scattered across long natural text. Our evaluations show that popular LLMs effectively utilize only 10-20% of the context and their performance declines sharply with increased reasoning complexity. Among alternatives to in-context reasoning, Retrieval-Augmented Generation methods achieve a modest 60% accuracy on single-fact question answering, independent of context length. Among context extension methods, the highest performance is demonstrated by recurrent memory transformers after fine-tuning, enabling the processing of lengths up to 50 million tokens. The BABILong benchmark is extendable to any length to support the evaluation of new upcoming models with increased capabilities, and we provide splits up to 10 million token lengths.
@@ 1 Introduction
Today, large language models (LLMs) and neural architectures are continually evolving and achieving remarkable improvements, particularly in their ability to handle longer contexts (OpenAI, 2023b; Reid et al., 2024; Anthropic, 2024). The ability of these models to process and generate text based on rich contextual information is crucial for several reasons. For example, longer contexts provide more information for the model to condition its outputs, leading to more accurate, contextually relevant, and up-to-date responses. Furthermore, long-context capabilities can enhance in-context learning by providing more in-context examples, instructions to follow, or example trajectories in context of reinforcement learning (Chevalier et al., 2023; Agarwal et al., 2024; Lee et al., 2024).
@@
Despite these advances in models capabilities, the benchmarks used to evaluate them have not kept pace. For example, current benchmarks such as Longbench (Bai et al., 2023) and L-Eval An et al. (2023) scale only up to 40,000 tokens, while models are capable of hundreds of thousands and millions of tokens (Rodkin et al., 2024; Reid et al., 2024; Bulatov et al., 2024; Anthropic, 2024; Liu et al., 2024a; Gu & Dao, 2023; OpenAI, 2023a).
@@
Figure 1: a) Generation of BABILong dataset. Facts relevant for the question are hidden inside a larger background texts from PG19. b) Recurrent transformers answer questions about facts from very long texts when retrieval augmented generation fails. Common RAG method fails to answer questions because order of facts matters. GPT-4 effectively uses only about 10% of the full 128K window. Gemini 1.5 Pro shows strong performance up to 64K tokens. Small LMs, ARMT & RMT with GPT-2 (137M) and Mamba (130M) fine-tuned for the task are able to solve it, with recurrent memory transformers scoring well up to record 50 000 000 tokens. Here we show the best results obtained by models.
@@
Creating natural and comprehensive long-context benchmarks that are human labeled is very challenging. As a consequence, synthetic benchmarks focusing on variations of "needle-in-a-haystack" tasks have become increasingly common (Zhang et al., 2024b; Liu et al., 2024a; Song et al., 2024b; Hsieh et al., 2024). One widely used needle-in-a-haystack task involves finding specific "needles with magic numbers" in a haystack of Paul Graham’s essays [^1] https://github.com/gkamradt/LLMTest_NeedleInAHaystack. However, the widespread use of this approach has highlighted its limitations - it is overly simplistic, and novel long context models often achieve perfect performance, as usually demonstrated by fully green heatmaps (Reid et al., 2024; Cohere, 2024; Liu et al., 2024a; Wang et al., 2024c). This shows that while it serves well as a basic verification tool, it is not a rigorous benchmark that can effectively challenge and differentiate advanced long-context models. Another major drawback of the original default setup1 is that model predictions are evaluated and scored by an LLM (GPT-3.5-turbo) on a scale of 1 to 10, with the same single needle used for each position and document length. While averaging over multiple different needles can provide more robust results.
@@
To bridge this gap, we introduce the BABILong benchmark, designed to test language models’ ability to reason across facts distributed in extremely long documents. BABILong includes a diverse set of 20 reasoning tasks, including fact chaining, simple induction, deduction, counting, and handling lists/sets, that were designed as prerequisites for any system that aims to be capable of conversing with a human (Weston et al., 2016). As a source of long natural documents we use books from PG19 corpora (Rae et al., 2020). In this way, BABILong allows the construction of tasks of almost arbitrary length, in order to adapt them to the evaluation of new, more powerful models in an extensible and controllable way. We provide sets of predefined lengths with splits up to 10 million tokens, and we evaluate models on samples with up to 50 million tokens.
@@
We find that popular LLMs effectively use only 10-20% of the context, with performance declining sharply as length and task complexity increase. Retrieval-Augmented Generation methods achieve a modest 60% accuracy in answering single-fact questions, regardless of context length. Among other methods, Mamba and Recurrent Memory Transformers (RMT and ARMT) show the highest performance, with ARMT capable of processing lengths up to 50 million tokens.
@@
2. We evaluate over 30 recent long-input language models with various sizes, architectures, and context extension methods on BABILong.
@@
4. We demonstrate successful in domain single fact question answering with the recurrent memory transformer on input texts up to 50 million tokens, which is a record for the sequence size processed by a single model.
The BABILong benchmark data and code for evaluation are available.222code: https://github.com/booydar/babilong, evaluation dataset: https://huggingface.co/datasets/RMT-team/babilong, leaderboard: https://huggingface.co/spaces/RMT-team/babilong
@@ 2 The BABILong Benchmark for Long Context Processing
The fundamental concept behind the Benchmark for Artificial Intelligence for Long-context evaluation is extending the length of existing tasks to test the ability of generative models in handling long contexts. Solving tasks with a long context size requires the model to distinguish important information from large amounts of irrelevant details. To simulate this behavior we "hide" the sentences of the original task between the sentences of irrelevant text that is drawn from another closely related distribution (see Figure 1a). Examples are constructed by gradually adding new sentences from the background dataset in their natural order until the augmented sample reaches the desired length. This way, we are not bound by the length of the original task itself, making it possible to assess even the longest available models with context sizes up to millions of tokens. For background text we use books from the PG19 dataset  (Rae et al., 2020) due to the substantial book lengths and naturally occurring long contexts. The model is required first to distinguish the sentences related to the original task, then memorize and subsequently utilize them to generate the correct solution.
@@ 3 Benchmarking Results
Table 2: BABILong is a challenging benchmark for current long-context models. Even models that claim to support 128K tokens experience degradation beyond 10% of their input capacity. RAG methods do not help, while fine-tuning of small scale models (ARMT and RMT, 137M and Mamba, 130M) shows that the tasks are solvable. Values represent average accuracy over QA1-QA5 tasks from BABILong. Models are grouped by the length of the context they claim to support.
To maximize value for the research community, we have included models with the highest number of monthly downloads [^3] As of May 2024. Since then we have added new models such as Mistral v0.3, LLama-3.1, Qwen-2.5, Phi-3.5, and GPT-4o-mini. from the Hugging Face platform in our evaluation such as LLama-3 (AI@Meta, 2024); 32k-64k – Mistral (Jiang et al., 2023), Mixtral (Jiang et al., 2024); 128k – ChatGLM3 (Du et al., 2022), LLama-3.1 (Dubey et al., 2024), Phi-3 and Phi-3.5 (Abdin et al., 2024), Command-R (Cohere, 2024), Qwen-2.5 (Team, 2024); 200k – Yi (Young et al., 2024); including long-context fine-tuning: LongChat (Li et al., 2023a), LLama-2-7b-32k [^4] https://huggingface.co/togethercomputer/Llama-2-7B-32K-Instruct, LongAlpaca (Chen et al., 2023); long-context adaptation methods: Yarnv2 Mistral (Peng et al., 2023b), Mistral and LLama-2 with Activation Beacons (Zhang et al., 2024a). As a reference, we included GPT-4 (gpt-4-0125-preview) and Gemini 1.5 Pro 002, currently the most powerful model available. Retrieval-augmented generation was also tested, as it represents a common solution for long document QA. As alternatives to traditional architectures, we considered the Mamba (Gu & Dao, 2023), Jamba (Lieber et al., 2024), Recurrent Memory Transformer (RMT) (Bulatov et al., 2022), and Associative RMT (ARMT) (Rodkin et al., 2024). Summary of evaluation results is presented in the Table 2. The table reports average accuracy of models on the first 5 tasks (QA1-QA5) of BABILong for different context sizes. For evaluation details for each task see Table 4 and Appendix C.
@@ 3.1 Evaluation of Effective Context Size
Our benchmarking results show that current LLMs do not efficiently use their full context (Fig. 2). Only 23 out of 34 tested LLMs were able to correctly answer 85% or more of the questions for any of QA1-QA3 tasks in a baseline setting without any background distractor text. Even for the the simplest task involving a single supporting fact (QA1), the majority of models are only able to efficiently use up to 4K tokens, except for GPT-4 and LLama-3.1-70b, which perform well up to 16K, as well as Qwen-2.5-70b and Gemini Pro 1.5 up to 64K. The range of full context utilization on QA1 varies from 5% to maximum 50%. When two supporting facts are required for an answer, only GPT-4 and Gemini Pro 1.5 can solve the task without background text. When facts are embedded within texts, all tested LLMs fall below 85% performance(Fig. 2, QA2). The task with three supporting facts proves to be extremely challenging to current LLMs, with the best scores falling below 80% (Fig. 2, QA3).
Going deeper in performance of specific models presented in the Table 2 we found the following. Yarn fails to extend to longer contexts despite showing stable results in long-context language modeling (Peng et al., 2023b). LongChat, LongAlpaca, and both LLama2-7B-32K and LLama2-7B-32K-instruct models, even when fine-tuned on 32K lengths, failed to perform well on 32K tokens. Activation Beacon performed better than Yarn context extension method for Mistral 7B, but still achieved low results (< 40%) on 32K contexts. In contrast, Mistral-v0.2 and Mixtral-v0.1, trained on lengths up to 32K, performed well on these lengths. Yi-9B-200k, trained on sequences up to 200K, shows less than 30% on 64K tokens and more. Yi-34B-200k shows very promising and stable results on lengths up to 64K, but unfortunately we were not able to run it on 128K tokens. Phi-3-mini drops significantly from 64K to 128K, reaching less than 10%, while Phi-3-medium maintains 30% at 128K. Jamba-v1 and Phi-3-mini show close results, but Jamba-v1 does not have drop at 128K and shows 34% on this length. Command-R and Phi-3-medium are the most robust to longer contexts, but start to lose performance more sharply at 128K. Phi-3-medium and Command-R show results very close to GPT-4 at 32K+ contexts.
We added results for models released since June 2024, including LLama 3.1, Phi 3.5, Qwen 2.5, and GPT-4o-mini. All claim to support 128K context length. Phi-3.5-mini shows improvements mainly for contexts up to 16K. The new Phi-3.5-MoE performs similarly to Phi-3-medium, but with only 6.6B active parameters compared to 14B in Phi-3-medium. LLama-3.1 models show significant improvement: LLama-3.1-8B matches GPT-4o-mini, while LLama-3.1-70B outperforms GPT-4 on longer contexts. Qwen-2.5 models outperform LLama-3.1 of similar size and achieve the best results of all evaluated open LLMs on BABILong.
Most of the new models use multistage pre-training with progressively increasing sequence lengths. For example, LLama-3.1 (Dubey et al., 2024) is pre-trained in six stages from 8K to 128K, and only proceeds to larger lengths if it maintains high short context scores and successfully solves the needle-in-a-haystack task. During supervised fine-tuning, LLama-3.1 models mix short context data with synthetic long context data, including question-answering, summarization, and code tasks.
@@ 3.3 Fine-Tuning Models on BABILong
We performed fine-tuning experiments with GPT-3.5-Turbo, Mistral-7B-Instruct-v0.2, RMT and ARMT with GPT-2 (137M) backbone, and Mamba (130M) models. Fine-tuning results are in Figure 3b and Appendix I Figure 9.
RMT with a GPT-2 (Radford et al., 2019) backbone model is trained on each task individually with a segment size of 512 and memory size of 16. ARMT with GPT-2 used 10 memory tokens (Rodkin et al., 2024). Train and evaluation splits of each task contain 10000 and 1000 samples, respectively, with a number of facts in each sample ranging from 2 to 320 depending on the task. A curriculum strategy is employed, initiating training on short sequences that fit in a single segment and then gradually introducing longer sequences once the training converges. During each curriculum stage $n$, sequence lengths are randomly sampled from $1$ to $n$ segments. We provide details of training and evaluation procedures in Appendix C.
RMT and ARMT models trained on 32 segments totalling in 16K tokens demonstrates strong performance on this length. Notably, recurrent memory models outperform GPT-4 significantly, underscoring the efficiency of memory mechanism in processing long context. Even more importantly, the power of recurrent models extends to sequences longer than the training size. RMT shows consistent performance on longer sequences, up to 128k tokens, with only a marginal quality degradation. Surprisingly, with context sizes scaling to 1 million, 10 million tokens, and even 11.1 million tokens, which is over 600 times of the training length. While ARMT successfully scales even further, reaching up to 50 million tokens (Rodkin et al., 2024). Recurrent memory models persistently outperform the larger counterparts utilizing RAG.
Finetuned recurrent models, Mamba, RMT and ARMT perform equally well on QA1, however due to the technical limitations of the Mamba implementation, the inference beyond 128k was extremely slow, which makes it nearly impossible to process longer sequences. Recurrent Memory models greatly outperform retrieval-augmented models and are able to process sequences up to 10M and 50M tokens much faster than Mamba. However, Mamba has an edge in complex tasks such as remembering a large number of facts in QA3 (Table 4).
We evaluated GPT-3.5 and Mistral-7B models fine-tuned with 1000 samples from QA1 for 3 epochs. The evaluation results are shown in Appendix I Figure 9. Fine-tuning dramatically improves performance for longer contexts making scores uniform across all input sizes. Still, these results are behind of fine-tuned Mamba, RMT and ARMT.